If Brad had $25.65 in quarters and dimes his money jar, but he had 20 more quarters than dimes, how many dimes did he have?
Answer by mananth(16949)


If Brad had $25.65 in quarters and dimes his money jar,
let number of dimes be x
Quarters will be x+20
10x+25(x+20)= 2565 (cents)
10x+25x+500=2565
35x = 2565-500
30x=2065
x=59 (dimes)
quarters = 59+20 =79
CHECK
79*25+59*10=2565